OpenStudy (anonymous):

how do you solve 1.4km = (b/2) ^ 1/3

OpenStudy (anonymous):

(14/10)^3=b/2 b=((14/10)^3)*2 write it like +((14/10)^3)*2 in an excel cell and get the result